HUBBLE SPACE TELESCOPE DAILY REPORT #849 PERIOD COVERED: 0800 EST Daily Status Report as of 082/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/BL 4222 (Mass Exchange Binaries- Cycle 1, Final Piece) The Faint Object Spectrograph (Blue Side) was used to make two spectral observations of the binary star system HZ Herculis. This system has a normal star as the central star, with a compact secondary companion. The first observations was received and displayed. It was reported to have an overlight sum 43,000. The second observation was lost because it completely filled the science tape recorder and the required down-link between the first and second observation was not available. As a result the science tape recorder was rewound to make room for the second observation, which was received and displayed. The OSS had data receipt problems, but was able to display the observation. It was reported to look okay. A line outage recorder playback was requested. 1.2 Completed WFC 1297 (GTO High Latitude Parallel Program) The Wide Field Camera was used to make two high latitude observations which were taken in parallel with the prime science observations. Both observations were received and displayed in the OSS. They were reported to look okay. 1.3 Completed WFC 4658 (Determination of the Extragalactic Distance Scale, M101) The Wide Field Camera was used to make two observations of galaxy M-101 which is located in Ursa Major. This is to attempt to discover new Cepheid variable stars. Both observations were executed in course track, with no apparent problems in the telemetry. The data was received and displayed, and both observations were reported to look okay. 1.4 Completed PC 3603 (Snapshot of Protoplanetary Nebulae) The Planetary Camera was used to attempt to make one observation of the protoplanetary system PK043+11D1. The observation was scheduled to be executed in fine lock. A bad V3 observer bias, left over from the high drift V3 drift rate prior just prior to the previous earth occultation, caused the guide star acquisition to fail fine lock. It did, however, obtain course lock. Just after the day/night transition, there was a big glitch in the V3 bias, but the lock held. Data was received and displayed in the OSS. The target was reported to be smeared as a result of the V3 jitter. 1.5 Completed FOC/96 4804 (UV Imaging of Nearby Galaxies With FOC/96) The Faint Object Camera was used to make one ten minute ultraviolet snapshot exposure of nearby galaxy UGC-5398 using filter F220W. The observations was executed in course track. A slow roll was observed in the V1 observer bias during the observation. The image still remains on the science tape recorder, and will be analyzed later. 1.6 Completed GHRS/2 4066 (GHRS Echelle Spectral Lamp Mini-Functional Test) The Echelle spectral lamp Mini-Fuction Test was completed as an internal test and calibration of the High Resolution Spectrograph. The first part of the test was received in real-time, and the second part was during LOS. The OSS expressed concerns about the lamp assembly temperature going out of limit, by being left on too long. It was monitored closely. 1.7 Completed PC 4140 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor Part 2) The Planetary Camera was used to make internal flat observations which will be used to look for new forms of contamination (i.e., daises and measles), and the return of quantum efficiency hysteresis. The observations complete as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 8 (Failure in FL at 081/2113Z, Bad Bias) (See Note 1.4) Scheduled Reacquisitions : 3 Successful : 3 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 22 Successful : 22 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete PC 3603 (Snapshots of Protoplanetary Nebulae) 3.2 Complete PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) 3.3 Complete FOC/96 4071 (UV Studies of Galaxies - Cycle 2) 3.4 Complete WFC 1297 (GTO High Latitude Parallel Program) 3.5 Complete GHRS/2 3737 (The Onset of Chromosheric Activity) 3.6 Complete S/C 4816 (FOC/48 Re-activation (Camera Section Only)) 3.7 Complete GHRS/2 1168 (Interstellar Carbon and Oxygen Cycle 2) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: FGS #2 Anomaly Investigation Further testing and analysis of the FGS #2 anomalous behavior are still ongoing. In the interim FGS #2 will not be used for observations. 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-